OpenStudy (angela210793):

Y=x*arctanx when does y'=0???

OpenStudy (lgbasallote):

i believe the first step would be to differentiate =_= this sounds complicated...

OpenStudy (anonymous):

is it y' =0 when x=0

OpenStudy (anonymous):

\[\frac{d(x*\tan^{-1}x)}{dx} = \frac{x}{1 + x^2} + \tan^{-1}x\]

OpenStudy (lgbasallote):

i wonder how we can solve for x \[\large \frac{x}{1+ x^2} + \tan^{-1} x = 0\]

OpenStudy (anonymous):

yup it is when x = 0 and no other solution

OpenStudy (lgbasallote):

we all know that..but *how* do we prove it..

OpenStudy (lgbasallote):

this is hard to prove i think..

OpenStudy (anonymous):

take tan(inverse)x on other side and draw graphs of both arctan(x) and -x/(1+x^2) you will find their graphs cutting each other at x = 0 only

OpenStudy (lgbasallote):

hmm i wonder if it's possible algebraically

OpenStudy (angela210793):

i have to study the monotony ,asimptotes and then graph it-_- and i haven't found the roots yet :(

OpenStudy (anonymous):

no not possible that way @igbasallote

OpenStudy (anonymous):

yes @angela210793 u have to study that to solve these type of questions.

OpenStudy (angela210793):

i know...but i can't find the roots :'(

OpenStudy (anonymous):

so better study these topics first then find the roots

OpenStudy (angela210793):

oh wait...(atctg)'=1/1+x^2 no?

OpenStudy (anonymous):

Yes you are doing wrong earlier..

OpenStudy (angela210793):

when y'=o?

OpenStudy (anonymous):

\[x + \tan^{-1}x^2 + \tan^{-1}x = 0\] \[x^2.\tan^{-1}x + x + \tan^{-1}x = 0\] \[D = b^2 - 4ac\] \[D = 1 - 4\tan^{-2}x\] \[x = \frac{-1 \pm \sqrt{1-4\tan^{-2}x}}{2\tan^{-1}x}\]

OpenStudy (anonymous):

Well, I am just trying.. Ha ha ha..

OpenStudy (angela210793):

:O i dont understand anyway can u check if this is right...two different methods gave me two different answers :(

OpenStudy (angela210793):